Solve the following boundary value problems.


(i) 𝑦


β€²β€² + 4𝑦 = 0; 𝑦(0) = 3, 𝑦(πœ‹/2) = βˆ’3,


(ii) 𝑦


β€²β€² βˆ’ 25𝑦 = 0; 𝑦(βˆ’2) = 𝑦(2) = cosh 10.


(iii) 𝑦


β€²β€² + 2𝑦


β€² + 2𝑦 = 0; 𝑦(0) = 1, 𝑦(πœ‹/2) = 0.


1
Expert's answer
2021-11-29T13:37:34-0500
  • The associated characteristic equation is

"\lambda^2+4=0" which gives the general solution "y= A\cos(2x)+B\sin(2x)". Now, applying the boundary conditions gives us : "A=3" and "B" arbitrary (as "\sin(0) =\sin(\pi)=0" ).

"y=3\cos(2x)+B\sin(2x)"

  • The associated characteristic equation is

"\lambda^2-25=0" which gives the general solution "Ae^{5x}+Be^{-5x}=A'\cosh(5x)+B'\sinh(5x)". Applying the boundary conditions yields "A'=1, B'=0", so "y=\cosh(5x)".

  • The associated characteristic equation is

"\lambda^2+2\lambda+2=0" which gives the general solution "e^{-x}(A\sin(x)+B\cos(x))". Applying the boundary conditions gives "B=1, A=0", so the solution is "y=e^{-x}\cos(x)"